### v1

@@ Abstract
Fragile web tests, primarily caused by locator breakages, are a persistent challenge in web development. Hence, researchers have proposed techniques for web-element re-identification in which algorithms utilize a range of element properties to relocate elements on updated versions of websites based on similarity scoring. In this paper, we replicate the original studies of the most recent propositions in the literature, namely the Similo algorithm and its successor, VON Similo. We also acknowledge and reconsider assumptions related to threats to validity in the original studies, which prompted additional analysis and the development of mitigation techniques. Our analysis revealed that VON Similo, despite its novel approach, tends to produce more false positives than Similo. We mitigated these issues through algorithmic refinements and optimization algorithms that enhance parameters and comparison methods across all Similo variants, improving the accuracy of Similo on its original benchmark by 5.62%. Moreover, we extend the replicated studies by proposing a larger evaluation benchmark (23$\times$ bigger than the original study) as well as a novel approach that combines the strengths of both Similo and VON Similo, called HybridSimilo. The combined approach achieved a gain comparable to the improved Similo alone. Results on the extended benchmark show that HybridSimilo locates 98.8% of elements with broken locators in realistic testing scenarios.
Keywords:
Web Testing; Locators; Test Robustness; Test Maintenance; Test Evolution
@@ 1 Introduction
Motivated by the will to understand the causes of these discrepancies and to address the aforementioned challenges, in this paper, we replicate the Similo similo and VON Similo vonsimilo studies, improving the experimental setting of the original papers to address the identified limitations and threats to the validity. More in detail: (1) we improve Similo by optimizing the attributes and weights of the original algorithm. We evaluated six new similarity functions to calculate the similarity between web element properties and optimize the weights assigned to these similarities using a genetic algorithm. We also analyze the capabilities of a novel hybrid version that combines the capabilities of Similo and VON Similo. (2) We collected a benchmark dataset of more than 10,000 element pairs from multiple websites and versions over the past five years. Our benchmark is 12$\times$ bigger than the original Similo benchmark and 23$\times$ bigger than the VON Similo benchmark, and it is more reflective of the fine-grained modifications occurring in real web sites. (3) We perform a fairer comparison between the original Similo and its extensions by evaluating all algorithms on the same benchmarks, both using the original as well as our new extended benchmark. While we were able to replicate the original study results, our findings are in contrast with ones by Nass et al. vonsimilo as we show that VON Similo underperforms Similo in directly identifying the target element but excels in identifying the visual overlap of the target element.
@@ 3.1 Similo
Similo requires a working version of the web application where the web element can be identified. Given the original web element $T$ and its properties $T.a_{1}...,T.a_{n}$, as well as all new web elements $C_{1},...,C_{n}$, of which one should be the original one $C^{\prime}$, Similo computes a similarity score between $T$ and each $C_{1},...,C_{n}$ individually:
@@ 3.2 VON Similo
The idea behind VON Similo is that web elements often consist of multiple parts. The DOM orders its elements so that a node’s children are inside the area the parent occupies on the visually rendered website. They appear as one visual unit to the user and can be interacted with as one unit.
@@
    The center of $E_{2}$ is located inside of $R_{1}$. (The paper describes this differently - “The center of the web element $W_{1}$ (here $E_{1}$) is contained in the rectangle $R_{1}$ but the underlying code shows that $E_{2}$ is compared with $R_{1}$).
@@
After calculating all overlapping elements $E_{1},...,E_{n}$ for one element $E$, we replace the properties $E.a_{1},...,E.a_{m}$ of that element with lists of property values of the overlapping nodes, meaning the properties of $E$ would look like this: $[E_{1}.a_{1},...,E_{n}.a_{1}],\allowbreak...,\allowbreak[E_{1}.a_{m},...,E_{n}.a_{m}]$.
@@ 3.5 Implementation
To be able to use the Similo algorithm in a practical setting, we implemented a library for Java, compatible with the Selenium WebDriver framework. The library wraps standard Selenium locators and automatically uses the Similo algorithm to identify the correct web element when the original locator fails. Upon first usage or locator changes, the library captures relevant element properties and saves them in an SQL database to do accurate matching in subsequent test executions. The wrapper supports all Selenium’s locator strategies (e.g., XPath, CSS selectors, ID), minimizing integration effort with existing test suites.
The implementation features built-in self-repair capabilities and intelligently updates locators within the database when elements change. Doing so, it enhances test resilience without modifying test code directly but remembering the connection between the locator and its current state of properties. Moreover, the library monitors Similo scores and issues configurable warnings when low score matches occur, guiding developers toward potentially broken locators. The library is available on GitHub and has comprehensive documentation replication-package .
@@ 4.3.1 Similo++
Compared Properties. Along with the properties used in the original Similo algorithm, we analyzed the frequency (how many element pairs have that property) and stability (in how many cases has the property the same value on the updated version). We found two additional suitable attributes: type (frequency of 12% and stability of 96%) and aria-label (frequency of 16% and stability of 82%). Additional we compare all attributes for that element as a key-value map.
@@
Jaro Winkler. The Jaro-Winkler distance is a string metric measuring the edit distance between two sequences. The Jaro-Winkler distance is given by a modification of the Jaro distance formula, where more weight is given to strings that match from the beginning. The Jaro distance is given by $d=1-\frac{1}{3}(\frac{m}{|s_{1}|}+\frac{m}{|s_{2}|}+\frac{m-t}{m})$, where $m$ is the number of matching characters and $t$ is half the number of transpositions. A transposition is a pair of matching characters in the wrong order in one of the strings. The Jaro-Winkler similarity is then calculated as Jaro $\text{Jaro similarity}+l\times p\times(1-\text{Jaro similarity})$, where $l$ is the length of the common prefix at the start of the string up to a maximum of four characters, and $p$ is a constant scaling factor, often 0.1. For example, the Jaro-Winkler similarity between “kitten” and “sitting” is approximately 0.74, assuming the Jaro distance is 0.77, and there is no common prefix.
@@
Exponential decay. uses the Euclidean distance but adds exponential decay, with differing $\lambda$ values calculated as $e^{-\lambda d}$, where $d$ is the Euclidean distance between the points. The benefit is that the function is already normalized to $[0,1]$ and approaches 00, eliminating the need to define a maximum distance. We used different values for $\lambda$, evaluating three exponential decay similarity functions, namely $\lambda=0.001$ as a small decay, $\lambda=0.005$ as a medium, and $\lambda=0.01$ as a large decay.
@@ 4.3.3 HybridSimilo
The HybridSimilo approach aims to overcome the limitations of VON Similo mentioned in T4. As discussed, approaches using visual overlap can only identify all elements in the visual overlap with the same accuracy, but not the exact element described by the locator. This can lead to problems in certain testing scenarios, as the element selected in the visual overlap might not be the exact target. On the other hand, Similo can identify the one single concrete element with high accuracy. A preliminary benchmark showed that the basic VON Similo algorithm could directly identify the target element in 85.5% of cases and rank it among the top five in 95.5% of cases. At the same time, Similo was able to identify the target element in 88% of cases directly and in 94% among the top five. While the basic VON Similo algorihm is superior in selecting all the elements in the visual overlap, it is inferior in selecting the exact target element.
The idea is to overcome the limitations of VON Similo by combining both alogrithms, leveraging the strenghts of both. We pre-select the elements in the visual overlap using VON Similo and then identify the concrete target among them using Similo. In theory, the visual overlap identified by VON Similo should contain elements which differ in their tags, attributes and properties, as the contribute to the formation of a visually cohesive unit. Similo should be able to correctly select the concrete target from the this pre-selection made by VON Similo. As the algorithm combines Similo and VON Similo, we named it HybridSimilo.
@@ 4.4 Metrics
Metric 5 $\cdot$ Locator Changed - Exact Match. This metric focuses on an algorithms performance specifically for elements where traditional unique locators (ID, XPath, or ID-XPath) have failed due to website updates. It is similar to Metric 4, which considers all element pairs, but Metric 5 focuses exclusively on the subset of elements that represent the cases where a real test suite would fail and manual locator repair would be needed.
By focusing on these cases, we can evaluate the effectiveness of different algorithms specifically in situations where conventional locator strategies fail between updates to the website. This metric provides a more focused assessment of the algorithm’s capability to address the main challenge in web test maintenance, correctly recovering from broken locators without manual intervention.
@@ 5 Results
Table 2 shows the results for RQ0-1-3.
The columns show the results for different metrics with the number of localization complying with the metric first and the percentage of the total number of localizations in brackets. The metric in parenthesis indicates the benchmark the algorithm was optimized on as well as the metrics being optimized by the genetic algorithm. I.e. (LLM, M4) means that the algorithm was optimized on the LLM VON Similo benchmark and the metric M4 was used as a fitness function. HybridSimilo is not applicable to M2, because the process of pre-selection does not work with the study setup, where a concrete element pair is given and the score needs be calculated. The columns are each for a specific variation of Similo. For the Similo algorithms taken from the original papers (Similo and VON Similo) are followed by the specific paper they are taken from in brackets. For the extended Similo algorithms (Similo++, VON Similo++, HybridSimilo) the content of the brackets indicates on which benchmark the algorithm was optimized and what metric was used as a fitness function. Results are reported separately for each considered benchmark. For metrics M2 and M6 the underlying numbers, i.e., the number of elements correctly classified by the threshold and the exact fitness have no informative value, why we only report the percentage.
@@ 5.1 Replication (RQ0)
For Similo, the replicated algorithm and benchmark found 88.99% elements, where the original paper was able to locate 88.64% of them. For VON Similo used in the LLM VON Similo paper, our replication found 91.65% of elements, where the original paper found 91.29% of elements. These minor differences were expected, as we reloaded the benchmarks from the Web Archive and differences in browser versions and window size can alter coordinates, shapes and areas as well as neighboring text used in the algorithm.
When replicating Similo and VON Similo algorithms on the VON Similo benchmark, we found challenges associated with the use of random non-fitting element pairs in the original benchmark, as the specific elements can change the results. Nonetheless, we found an accuracy of 93.64% for VON Similo, compared to 94.1% in the original paper, and a replicated accuracy of 85.22% for Similo compared to the original 82.30%. Based on these results, we consider our reproduction of the original results to be successful.
@@ 5.5.2 External Validity
The limited number of websites in our evaluation poses a threat in terms of generalizability of our results to other web apps. We assume that the datasets comprising the most popular websites provides a realistic representation of how websites change over time, as seen in a continuous integration environment. Furthermore, only the static front pages of websites were used for the dataset. Front pages typically consist of links, headers, images, and menu items and may not represent the diversity of elements found in other parts of a web application. This selection bias could affect the algorithm’s effectiveness on pages different than front pages, where elements such as selects, tables, and table items appear more frequently. The final algorithm, with its optimized weights and similarity functions, is tailored to this dataset. Despite using cross-validation to prevent overfitting, there is a risk that the results are overly optimized for the dataset, potentially affecting the algorithm’s performance on different websites. The data was scraped from the Wayback Machine, which may not always capture a website’s complete or accurate representation. Additionally, the snapshots are scraped across different browsers in different countries and further pre-processed before saving and rendering them, introducing further inaccuracies into the dataset. All these factors could lead the optimized algorithm to perform differently in sanitized testing environments.
@@ 6 Discussion
Discussing Main Differences. Column M2 (i.e. the metric used for the VON Similo algorithm) differs significantly from the other metrics, as it is the only metric on which VON Similo outperforms all other algorithms. This difference is primarily due to variations in the study setup compared to those used for Similo and LLM VON Similo. The VON Similo reported than 94.1% of element pairs were accurately classified as either matching or non-matching. Given that in case of a localization in a test case the target $T$ will be compared with all $C$, we would require $|C|-1$ correct classifications as non-matching (sensitivity being 0.97) and one as matching (recall being 0.922). With approximately 800 elements in $C$ to compare with $T$, the probability of correctly identifying the target in the new version is calculated as $0.97^{799}\cdot 0.92$, which approximates to 00.
@@
Impact of long and short inter-version time intervals on performance and benchmark. At first glance, it may appear that improved methods offer negligible benefits for benchmarks with shorter inter-version intervals since baseline algorithms already perform strongly. However, considering improvements relative to the maximum achievable performance reveals their true significance. For instance, in long-term intervals, performance improved from $86.6\%$ to $91.7\%$, representing $38\%$ of the possible remaining margin towards $100\%$. For short-term intervals, the baseline already achieved $99.0\%$, and improved to $99.7\%$, again representing a substantial $70\%$ of remaining possible improvement. Furthermore, focusing specifically on elements with changed locators, performance improved from $95.8\%$ to $98.8\%$, covering an even larger relative improvement of $71\%$. These results underline that even minor absolute improvements at high baseline levels can translate into significant practical gains, particularly when addressing more challenging locator updates.

### v2

@@ Abstract
[…]
@@ 1 Introduction
[…]
@@ 3.1 Similo
Algorithm 1 Similo
1:function Similo(target_element,candidate_elementstarget\_element,candidate\_elements)
2:  $best\_score\leftarrow 0$
3:  $best\_candidate\leftarrow\text{null}$
4:  for all $candidate\in candidate\_elements$ do
5:   $score\leftarrow$ SimiloScore(target_element,candidatetarget\_element,candidate)
6:   if score>best_scorescore>best\_score then
7:    $best\_score\leftarrow score$
8:    $best\_candidate\leftarrow candidate$
9:   end if
10:  end for
11:  return best_candidatebest\_candidate
12:end function
13:function SimiloScore($target,candidate$)
14:  $total\leftarrow 0$
15:  for all $property\in target.properties$ do
16:   $t\leftarrow$ Get($target,property$)
17:   $c\leftarrow$ Get($candidate,property$)
18:   $sim\leftarrow$ Similarity($t,c$)
19:   $total\leftarrow total+sim\times WEIGHTS[property]$
20:  end for
21:  return totaltotal
22:end function
Similo (Algorithm 1) requires a working version of the web application where the web element can be identified. Given the original web element $T$ and its properties $T.a_{1}...,T.a_{n}$, as well as all new web elements $C_{1},...,C_{n}$, of which one should be the original one $C^{\prime}$, Similo computes a similarity score between $T$ and each $C_{1},...,C_{n}$ individually:
@@ 3.2 VON Similo
Algorithm 2 Von Similo
1:original element $T$, old web page OldSiteOldSite, candidate elements $C$
2:best matching element in $C$
3:function VonSimilo($T,OldSite,C$)
4:  $T\_overlap\leftarrow$ Overlap($T,OldSite$)
5:  UpdateProperties(T,T_overlapT,T\_overlap)
6:  $best\_score\leftarrow 0$
7:  $best\_candidate\leftarrow\text{null}$
8:  for all $cand\in C$ do
9:   $cand\_overlap\leftarrow$ Overlap($cand,C$)
10:   UpdateProperties(cand,cand_overlapcand,cand\_overlap)
11:   $score\leftarrow$ VonSimiloScore($T,cand$)
12:   if score>best_scorescore>best\_score then
13:    $best\_score\leftarrow score$
14:    $best\_candidate\leftarrow cand$
15:   end if
16:  end for
17:  return best_candidatebest\_candidate
18:end function
19:function VonSimiloScore($T,cand$)
20:  $total\leftarrow 0$
21:  for all property lists $(L_{T},L_{c})$ in Zip($T.properties,cand.properties$) do
22:   $m\leftarrow 0$
23:   for all $p_{T}\in L_{T}$ do
24:    for all $p_{c}\in L_{c}$ do
25:      $s\leftarrow$ Similarity($p_{T},p_{c}$)
26:      if $s>m$ then
27:        $m\leftarrow s$
28:      end if
29:    end for
30:   end for
31:   $total\leftarrow total+m$
32:  end for
33:  return totaltotal
34:end function
35:function UpdateProperties($e,OverlapSet$)
36:  for all $prop\in e.properties$ do
37:   $values\leftarrow[\ ]$
38:   for all $o\in OverlapSet$ do
39:    append $o.properties[prop.name].value$ to valuesvalues
40:   end for
41:   $prop.value\leftarrow values$
42:  end for
43:end function
44:function Overlap($e,S$)
45:  $O\leftarrow[\ ]$
46:  for all $cand\in S$ do
47:   if Overlaps($e,cand$) then
48:    append candcand to $O$
49:   end if
50:  end for
51:  return $O$
52:end function
53:function Overlaps($e,cand$)
54:  $r_{T}\leftarrow e.rect()$
55:  $r_{C}\leftarrow cand.rect()$
56:  $ratio\leftarrow\textsc{IntersectArea}(r_{T},r_{C})\,/\,\textsc{UnionArea}(r_{T},r_{C})$
57:  $inside\leftarrow\textsc{IsInside}(cand.center(),r_{T})$
58:  return $(inside\wedge ratio\geq 0.85)$
59:end function
The idea behind VON Similo (Algorithm 2) is that web elements often consist of multiple parts. The DOM orders its elements so that a node’s children are inside the area the parent occupies on the visually rendered website. They appear as one visual unit to the user and can be interacted with as one unit.
@@
    The center of $E_{2}$ is located inside of $R_{1}$. (The paper describes this differently - “The center of the web element $W_{1}$ (here $E_{1}$) is contained in the rectangle $R_{1}$ - but the underlying code shows that $E_{2}$ is compared with $R_{1}$).
@@
After calculating all overlapping elements $E_{1},...,E_{n}$ for one element $E$, we replace the properties $E.a_{1},...,E.a_{m}$ of that element with lists of property values of the overlapping nodes, meaning the properties of $E$ would look like this: $E_{1}.a_{1},...,E_{n}.a_{1}],\allowbreak...,\allowbreak[E_{1}.a_{m},...,E_{n}.a_{m}]$.
@@ 3.5 Implementation
To be able to evaluate the Similo algorithm in a practical setting, we implemented a library for Java, compatible with the Selenium WebDriver framework. The library wraps standard Selenium locators and automatically uses the Similo algorithm to identify the correct web element when the original locator fails. Upon first usage or locator changes, the library captures relevant element properties and saves them in an SQL database to do accurate matching in subsequent test executions. The wrapper supports all Selenium’s locator strategies (e.g., XPath, CSS selectors, ID), minimizing integration effort with existing test suites.
The implementation features built-in self-repair capabilities by locating the element whose locator is broken and using it to continue test execution. It also intelligently updates locators within the database when elements change by determining the most stable of the basic locators XPath, ID and ID-XPath and using it for future localizations. Doing so, it enhances test resilience without modifying test code directly but remembering the connection between the locator and its current state of properties. Moreover, the library monitors Similo scores and issues configurable warnings when low score matches occur, guiding developers toward potentially broken locators. The library is available on GitHub and has comprehensive documentation replication-package.
@@ 4.3.1 Similo++
Compared Properties. Along with the properties used in the original Similo algorithm, we analyzed the frequency (how many element pairs have that property) and stability (in how many cases has the property the same value on the updated version). We found two additional suitable attributes: type (frequency of 6-7% and stability of 95-96%) and aria-label (frequency of 10-12% and stability of 81-84%) across five random subsets of the training data as well as the cross validation set. Additional we compare all attributes for that element as a key-value map.
@@
Jaro Winkler. The Jaro-Winkler distance is a string metric measuring the edit distance between two sequences. The Jaro-Winkler distance is given by a modification of the Jaro distance formula, where more weight is given to strings that match from the beginning. The Jaro distance is given by $d=1-\frac{1}{3}(\frac{m}{|s_{1}|}+\frac{m}{|s_{2}|}+\frac{m-t}{m})$, where $m$ is the number of matching characters and $t$ is half the number of transpositions. A transposition is a pair of matching characters in the wrong order in one of the strings. The Jaro-Winkler similarity is then calculated as $\text{Jaro similarity}+l\times p\times(1-\text{Jaro similarity})$, where $l$ is the length of the common prefix at the start of the string up to a maximum of four characters, and $p$ is a constant scaling factor, often 0.1. For example, the Jaro-Winkler similarity between “kitten” and “sitting” is approximately 0.74, assuming the Jaro distance is 0.77, and there is no common prefix.
@@
Exponential decay. uses the Euclidean distance but adds exponential decay, with differing $\lambda$ values calculated as $e^{-\lambda d}$, where $d$ is the Euclidean distance between the points. The benefit is that the function is already normalized to $[0,1]$ and approaches 0, eliminating the need to define a maximum distance. We used different values for $\lambda$, evaluating three exponential decay similarity functions, namely $\lambda=0.001$ as a small decay, $\lambda=0.005$ as a medium, and $\lambda=0.01$ as a large decay.
@@ 4.3.3 HybridSimilo
The HybridSimilo approach aims to overcome the limitations of VON Similo mentioned in T4. As discussed, approaches using visual overlap can only identify all elements in the visual overlap with the same accuracy, but not the exact element described by the locator. This can lead to problems in certain testing scenarios, as the element selected in the visual overlap might not be the exact target. On the other hand, Similo can identify the one single concrete element with high accuracy. A preliminary benchmark showed that the basic VON Similo algorithm could directly identify the target element in 85.5% of cases and rank it among the top five in 95.5% of cases. At the same time, Similo was able to identify the target element in 88% of cases directly and in 94% among the top five. While the basic VON Similo algorithm is superior in selecting all the elements in the visual overlap, it is inferior in selecting the exact target element.
The idea is to overcome the limitations of VON Similo by combining both algorithms, leveraging the strengths of both. We pre-select the elements in the visual overlap using VON Similo and then identify the concrete target among them using Similo. In theory, the visual overlap identified by VON Similo should contain elements which differ in their tags, attributes and properties, as the contribute to the formation of a visually cohesive unit. Similo should be able to correctly select the concrete target from the this pre-selection made by VON Similo. As the algorithm combines Similo and VON Similo, we named it HybridSimilo.
@@ 4.4 Metrics
Metric 5 $\cdot$ Locator Changed - Exact Match. This metric focuses on an algorithms performance specifically for elements where traditional unique locators (ID, XPath, or ID-XPath) have failed due to website updates. It is similar to Metric 4, which considers all element pairs, but Metric 5 focuses exclusively on the subset of elements that represent the cases where a real test suite would fail and manual locator repair would be needed. By focusing on these cases, we can evaluate the effectiveness of different algorithms specifically in situations where conventional locator strategies fail between updates to the website. This metric provides a more focused assessment of the algorithm’s capability to address the main challenge in web test maintenance, correctly recovering from broken locators without manual intervention.
@@ 5 Results
Table 2 shows the results for RQ0-1-3. The columns show the results for different metrics with the number of localization complying with the metric first and the percentage of the total number of localizations in brackets. The metric in parenthesis indicates the benchmark the algorithm was optimized on as well as the metrics being optimized by the genetic algorithm. I.e. (LLM, M4) means that the algorithm was optimized on the LLM VON Similo benchmark and the metric M4 was used as a fitness function. HybridSimilo is not applicable to M2, because the process of pre-selection does not work with the study setup, where a concrete element pair is given and the score needs be calculated. The columns are each for a specific variation of Similo. For the Similo algorithms taken from the original papers (Similo and VON Similo) are followed by the specific paper they are taken from in brackets. For the extended algorithms (Similo++, VON Similo++, HybridSimilo) the content of the brackets indicates on which benchmark the algorithm was optimized and what metric was used as a fitness function. Results are reported separately for each considered benchmark. For metrics M2 and M6 the underlying numbers, i.e., the number of elements correctly classified by the threshold and the exact fitness have no informative value, why we only report the percentage.
@@ 5.1 Replication (RQ0)
For Similo, the replicated algorithm and benchmark found 88.99% elements, where the original paper was able to locate 88.64% of them. For VON Similo used in the LLM VON Similo paper, our replication found 91.65% of elements, where the original paper found 91.29% of elements. These minor differences were expected, as we reloaded the benchmarks from the Web Archive and differences in browser versions and window size can alter coordinates, shapes and areas as well as neighboring text used in the algorithm. When replicating Similo and VON Similo algorithms on the VON Similo benchmark, we also found challenges associated with the use of random non-fitting element pairs in the original benchmark, as the specific elements can change the results.
RQ0 (replication): Our reproduced results closely match the original outcomes for both Similo and VON Similo, indicating that our replication was successful.
@@ 5.2 Comparison (RQ1)
RQ1 (comparison): Similo is more effective than VON Similo at retrieving the exact target element, while VON Similo performs better when broader visual or textual overlap is sufficient.
@@ 5.3 Improvements (RQ2)
RQ2 (improvements): Optimizing the property set, similarity functions, and weights yields clear improvements for both Similo and VON Similo, showing that their default configurations are not optimal. Key attributes (e.g., name, type, aria-label, text, location) consistently receive high importance, while others contribute little. Gains are likely even larger when optimizing for specific websites or frameworks.
@@ 5.4 Hybrid (RQ3)
To ensure that we did not overfit on the benchmark data, we conducted temporal cross-validation on the extended benchmark, using the last $n$ snapshots over all sites as test data and training on all earlier snapshots in five folds. This setup best reflects the common use case of the algorithm, where it is deployed in a functioning test suite and later tasked with locating changed elements on the same website. This approach ensures that our results translate to common testing scenarios. The results show that the performance of the optimized versions on the extended benchmark differs by only 0.1% on all metrics when using cross-validation. Similo++ and HybridSimilo again performing similarly.
RQ3 (hybrid): Hybridizing Similo and VON Similo offers benefits only in specific cases. VON Similo is useful for generating an initial shortlist when versions differ substantially, and Similo can then refine this selection to identify the exact element. This hybrid approach improves accuracy on benchmarks with large visual or structural changes, but provides little advantage when updates are minor. Temporal cross-validation confirms that these effects generalize, with hybrid and optimized Similo performing similarly.
@@ 5.5.2 External Validity
The limited number of websites in our evaluation poses a threat in terms of generalizability of our results to other web apps. We assume that the datasets comprising the most popular websites provides a realistic representation of how websites change over time, as seen in a continuous integration environment. However, different web frameworks and libraries that are not represented in our dataset could yield different results, and systematically evaluating them is left for future research. Furthermore, only the static front pages of websites were used for the dataset. Front pages typically consist of links, headers, images, and menu items and may not represent the diversity of elements found in other parts of a web application. This selection bias could affect the algorithm’s effectiveness on pages different than front pages, where elements such as selects, tables, and table items appear more frequently. The final algorithm, with its optimized weights and similarity functions, is tailored to this dataset. Despite using cross-validation to prevent overfitting, there is a risk that the results are overly optimized for the dataset, potentially affecting the algorithm’s performance on different websites. The data was scraped from the Wayback Machine, which may not always capture a website’s complete or accurate representation. Additionally, the snapshots are scraped across different browsers in different countries and further pre-processed before saving and rendering them, introducing further inaccuracies into the dataset.
All these factors could lead the optimized algorithm to perform differently in sanitized testing environments.
@@ 6 Discussion
Discussing Main Differences. Our results confirm that our Similo and VON Similo replications closely match the original studies and Similo consistently outperforms VON Similo on localizing a concrete element, while VON Similo is only superior on broader overlap metrics (M2/M3). When optimizing properties, similarity functions, and weights yields small but meaningful gains at high baselines (e.g., Similo++ and VON Similo++ improve exact-match accuracy by up to 5.6 percentage points and achieve up to 71% relative improvement on broken-locator cases) and a hybrid variant using VON Similo++ for candidate pre-selection and Similo++ for final ranking delivers the best overall performance on the original benchmarks and remains competitive on the extended dataset. This indicates that VON Similo is most useful as a overlap pre-filter under larger changes.
Column M2 (i.e. the metric used for the VON Similo algorithm) differs significantly from the other metrics, as it is the only metric on which VON Similo outperforms all other algorithms. This difference is primarily due to variations in the study setup compared to those used for Similo and LLM VON Similo. The VON Similo reported than 94.1% of element pairs were accurately classified as either matching or non-matching. Given that in case of a localization in a test case the target $T$ will be compared with all $C$, we would require |C|1|C|-1 correct classifications as non-matching (sensitivity being 0.97) and one as matching (recall being 0.922). With approximately 800 elements in $C$ to compare with $T$, the probability of correctly identifying the target in the new version is calculated as $0.97^{799}\cdot 0.92$, which approximates to 0.
@@
Impact of long and short inter-version time intervals on performance and benchmark. At first glance, it may appear that improved methods offer negligible benefits for benchmarks with shorter inter-version intervals since baseline algorithms already perform strongly. However, considering improvements relative to the maximum achievable performance reveals their true significance. For instance, in long-term intervals, performance improved from $86.6\%$ to $91.7\%$, representing $38\%$ of the possible remaining margin towards $100\%$. For short-term intervals, the baseline already achieved $99.0\%$, and improved to $99.7\%$, again representing a substantial $70\%$ of remaining possible improvement. Furthermore, focusing specifically on elements with changed locators, performance improved from $95.8\%$ to $98.8\%$, covering an even larger relative improvement of $71\%$. These results underline that even minor absolute improvements at high baseline levels can translate into significant practical gains, particularly when addressing more challenging locator updates. Additionally, we can see that while the Similo algorithms perform better on short-term changes, repairing locators with 98.8% accuracy, their performance on long-term changes of 1–5 years is significantly lower at 88.4%, though still better than baseline approaches. This underlines the need for frequent updates to the attribute set of a locator, as done in our implementation.